# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
    For any $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}\in S$, further let $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, where $\lambda_{i}^{x},\mu_{i}^{x}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$, it holds that
@@
    For any $(n-1)$-dimensional $L^{0}$-simplex $S_{1}$ contained in $S$ and any $L^{0}$-simplicial subdivision $\mathscr{S}$ of $S_{1}$, $\phi|_{ext_{L^{0}}(\mathscr{S})}$ is a proper $L^{0}$-labeling function of $\mathscr{S}$.
@@
Since for any $x\in S$ there exists an unique $\{\lambda_{i}^{x},i=1\sim n\}\subseteq L^{0}(\mathcal{F},[0,1])$ such that $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=1$, where $(\lambda_{1}^{x},\lambda_{2}^{x},\cdots,\lambda_{n}^{x})$ is called the $L^{0}$-barycentric coordinate of $x$. For brevity and clarity, in the proof, we always use $\lambda_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $x$ and $\mu_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $f(x)$. Then, for any $x\in S$, $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, we claim that
@@
namely
$a:=[(\lambda_{j}^{x}>0)\wedge(\lambda_{j}^{x}<\mu_{j}^{x})]\wedge\{\bigwedge_{i\in\{1,2,\cdots,n\}\backslash\{j\}}[(\lambda_{i}^{x}=0)\vee(\lambda_{i}^{x}<\mu_{i}^{x})]\}>0.$
@@
which contradicts with $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$.
@@
(1). For any $x\in S$, it is clear that $\phi(x)\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ and
@@
(2). For any $\{c_{k},k\in\mathbb{N}\}\in p(1)$ and any sequence $\{y_{k},k\in\mathbb{N}\}$ in $ext_{L^{0}}(\mathscr{S})$, since $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{y_{k}}x_{i}$ and $f(y_{k})=\sum_{i=1}^{n}\mu_{i}^{y_{k}}x_{i}$ for each $k\in\mathbb{N}$, we have
@@
for any $i\in\{1,2,\cdots,n\}$. Then, we have
@@
which implies that $\phi$ is $\sigma$-stable, namely $\phi$ is an $L^{0}$-labeling mapping of $\mathscr{S}$.
@@
for any $i\in\{1,2,\cdots,n\}$. Thus, $\phi$ is proper. ∎
Lemma 4.9 below is the $L^{0}$-simplex version of random Brouwer fixed point theorem in $RN$ modules.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $RN$ module $E$ and $f:S\rightarrow S$ be a $\sigma$-stable and random sequentially continuous mapping. Then $f$ has a fixed point.
@@
Let $\mathscr{S}$ be an $L^{0}$-barycentric subdivision of $S$. For $\phi$ and $\mathscr{S}$, by Theorem 2.9 there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{1}:=Conv_{L^{0}}(\{y_{1}^{1},y_{2}^{1},\cdots,y_{n}^{1}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}$, namely there exist a finite family $\{S_{k}^{1},k=1\sim M\}$ of $(n-1)$-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{k}^{1},k=1\sim M\}\in p(1)$ such that
@@
Let $\mathscr{S}_{1}$ be an $L^{0}$-barycentric subdivision of $S^{1}$. Now, similarly, for $\phi$ and $\mathscr{S}_{1}$, by Theorem 2.9 there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{2}:=Conv_{L^{0}}(\{y_{1}^{2},y_{2}^{2},\cdots,y_{n}^{2}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{1}$, such that $S^{2}$ is completely labeled by $\phi$ and satisfies that
@@
By induction, for each $m\in\mathbb{N}$, there exists an $(n-1)$-dimensional $L^{0}$-simplex $S^{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{m-1}$, such that $S^{m}$ is completely labeled by $\phi$ and satisfies that
@@
Then $\{(y_{1}^{m},y_{2}^{m}\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ is a sequence in $S^{(n)}$, where $S^{(n)}$ is the $n$-th Cartesian power of $S$.
For $\{y_{1}^{m},m\in\mathbb{N}\}$, since $S$ is random sequentially compact by Remark 4.4, there exists a random subsequence $\{y_{1}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to some $y$ in $S$. Since $\{diam(S^{m}),m\in\mathbb{N}\}$ converges $a.s.$ (of course, also in $\mathcal{T}_{\varepsilon,\lambda}$) to 00, by Lemma 3.5 of [26] $\{d^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to 00, where
@@
which implies that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$. To sum up, $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}})$, $k\in\mathbb{N}\}$ is a random subsequence of $\{(y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ for any $i\in\{1,2,\cdots,n\}$. Since $f$ is random sequentially continuous, by Lemma 4.4 of [26] there exists a random subsequence $\{(y_{1}^{m_{k}},y_{2}^{m_{k}},\cdots,y_{n}^{m_{k}}),k\in\mathbb{N}\}$ of $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}}),k\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ and $\{f(y_{i}^{m_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(y)$ for each $i\in\{1,2,\cdots,n\}$.
Let $y=\sum_{i=1}^{n}\lambda_{i}x_{i}$ and $f(y)=\sum_{i=1}^{n}\mu_{i}x_{i}$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ are such that $\sum_{i=1}^{n}\lambda_{i}=\sum_{i=1}^{n}\mu_{i}=1$. Next, we will show that $f(y)=y$, namely $y$ is a fixed point of $f$.
@@
$=j,$
then $\lambda_{j}^{m_{k},j}\geq\mu_{j}^{m_{k},j}$ by Lemma 4.8.
Furthermore, by Lemma 5.5 of [26], $\{\lambda_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\lambda_{j}$ and $\{\mu_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\mu_{j}$ for each $j\in\{1,2,\cdots,n\}$, then
@@
Lemma 4.9 is a generalization of [11, Theorem 2.3], since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is an $RN$ module, and a $\sigma$-stable a.s. sequentially continuous mapping is, obviously, random sequentially continuous, as noted in Remark 4.6. Since random sequential continuity is much more complicated than a.s. sequential continuity, Lemma 4.4 of [26] has played a key role in the treatment of random sequential continuity in the proof of Lemma 4.9. Besides, we would also like to emphasize another difference between the proofs of Lemma 4.9 and [11, Theorem 2.3]: the sequence $\{S^{m},m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes, which is constructed in Lemma 4.9, has the nice property that $S^{m+1}\subseteq S^{m}$ for each $m\in\mathbb{N}$, but the sequence $\{S^{m},m\in\mathbb{N}\}$ constructed in [11, Theorem 2.3] does not have this property. The nice property of our sequence will be used to study other problems in the future. Besides, the proof of Lemma 4.9 is simpler and clearer than that of [11, Theorem 2.3] since we only employ a single proper $L^{0}$-labeling function throughout the proof of Lemma 4.9.
Let $H$ be a Hilbert space and $G$ be a nonempty closed convex subset of $H$, then it is well known that the classical projection $P_{G}$ of $H$ onto $G$ is nonexpensive. By a completely similar method, one can prove the following lemma:
@@
Let $(H,\langle\cdot,\cdot\rangle)$ be a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ and $G$ be a nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $H$. Then, for any $x\in H$, there exists an unique element $P_{G}(x)$ in $G$ such that
@@
[…]
@@
In Lemma 4.11, since $G$ is $\sigma$-stable and $P_{G}$ is $L^{0}$-Lipschitzian, then $P_{G}$ is $\sigma$-stable by Lemma 2.11 of [27].
@@
Since $G$ is $a.s.$ bounded, there exists an $L^{0}$-simplex $S$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ such that $G\subseteq S$. Since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module, let $h:S\rightarrow G$ be the restriction of $P_{G}$ to $S$, where $P_{G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$, it is clear that $h$ is $\mathcal{T}_{c}$-continuous.
@@
Proposition 4.12 below was first established as Lemma 4.7 of [26], which is merely a reformulation of Theorem 2.1 in the case of a finitely generated $RN$ module. For the proof of Proposition 4.12, we refer the reader to [26].
@@
5 The equivalence of Proposition 2.3 and Proposition 2.4 together with other interesting results
@@
Sufficiency. For any $a\in B_{\mathcal{F}}$ and any $x,y$ in $G$, from the $\sigma$-stability of $h$ one has that $h(I_{a}x+I_{a^{c}}y)=I_{a}h(x)+I_{a^{c}}h(y)$. Then, when $I_{a}x=I_{a}y$, $h(y)=I_{a}h(x)+I_{a^{c}}h(y)$, so $I_{a}h(x)=I_{a}h(y)$, which amounts to saying that $h$ is local. ∎
Lemma 5.2 below is a special case of the main result of [42], which is enough for our purpose.
@@
Let $X$ and $Y$ be two complete separable metric spaces, $(\Omega,\mathcal{F},P)$ a complete probability space and $h:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ a local and continuous-in-probability operator. Then there exists a Carathéodory function $f:\Omega\times X\rightarrow Y$ (namely $f(\cdot,x)$ is measurable for any $x\in X$ and $f(\omega,\cdot)$ is continuous for any $\omega\in\Omega$) such that $h=h_{f}$. Here, $h_{f}:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ is defined by $h_{f}(x)=$ the equivalence class of $f(\cdot,x^{0}(\cdot))$, where $x^{0}$ is an arbitrarily chosen representative of $x$.
In fact, Lemma 5.2 has implied that $h$ is a.s. sequentially continuous, namely $\{h(x_{n}),n\in\mathbb{N}\}$ converges a.s. to $h(x)$ whenever $\{x_{n},n\in\mathbb{N}\}$ converges a.s. to $x$.

### v5

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
    For any $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}\in S$, further let $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, where $\lambda_{i}^{x},\mu_{i}^{x}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$. Then we have
@@
    For any (n1)(n-1)-dimensional $L^{0}$-simplex $S_{1}$ contained in $S$ and any $L^{0}$-simplicial subdivision $\mathscr{S}$ of $S_{1}$, $\phi|_{ext_{L^{0}}(\mathscr{S})}$ is a proper $L^{0}$-labeling function of $\mathscr{S}$.
@@
Since for any $x\in S$ there exists a unique $\{\lambda_{i}^{x},i=1\sim n\}\subseteq L^{0}(\mathcal{F},[0,1])$ such that $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $\sum_{i=1}^{n}\lambda_{i}^{x}=1$, where $(\lambda_{1}^{x},\lambda_{2}^{x},\cdots,\lambda_{n}^{x})$ is called the $L^{0}$-barycentric coordinate of $x$. For brevity and clarity, in the proof, we always use $\lambda_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $x$ and $\mu_{i}^{x}$ for the $i$-th $L^{0}$-barycentric coordinate of $f(x)$. Then, for any $x\in S$, $x=\sum_{i=1}^{n}\lambda_{i}^{x}x_{i}$ and $f(x)=\sum_{i=1}^{n}\mu_{i}^{x}x_{i}$, we claim that
@@
namely,
[…]
@@
contradicting the facts that $\sum_{i=1}^{n}\lambda_{i}^{x}=\sum_{i=1}^{n}\mu_{i}^{x}=1$.
@@
(1) For any $x\in S$, it is clear that $\phi(x)\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ and
@@
(2) For any $\{c_{k},k\in\mathbb{N}\}\in p(1)$ and any sequence $\{y_{k},k\in\mathbb{N}\}$ in $ext_{L^{0}}(\mathscr{S})$, since $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{y_{k}}x_{i}$ and $f(y_{k})=\sum_{i=1}^{n}\mu_{i}^{y_{k}}x_{i}$ for each $k\in\mathbb{N}$, we have
@@
for any $i\in\{1,2,\cdots,n\}$. Thus, we have
@@
which implies that $\phi$ is $\sigma$-stable, namely, $\phi$ is an $L^{0}$-labeling mapping of $\mathscr{S}$.
@@
for any $i\in\{1,2,\cdots,n\}$. Hence $\phi$ is proper. ∎
Proposition 4.9 below will be used in the proof of Lemma 4.10.
Proposition 4.9 ([53, Proposition 2.14]).
Let $(E_{1},\|\cdot\|)$ and $(E_{2},\|\cdot\|)$ be two $RN$ modules over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$, $G_{1}\subset E_{1}$ and $G_{2}\subset E_{2}$ two $\sigma$-stable subsets, and $f:G_{1}\rightarrow G_{2}$ a $\sigma$-stable random sequentially continuous mapping. For a sequence $\{(x_{1}^{m},x_{2}^{m},\cdots,x_{l}^{m}),m\in\mathbb{N}\}$ in $G_{1}^{l}$, where $l$ is a fixed positive integer and $G^{l}_{1}$ is the ll–th Cartesian power set of $G_{1}$, if there exists a random subsequence $\{(x_{1}^{M_{n}^{(0)}},x_{2}^{M_{n}^{(0)}},\cdots,x_{l}^{M_{n}^{(0)}}),n\in\mathbb{N}\}$ of which such that $\{x_{i}^{M_{n}^{(0)}},n\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to some $y_{i}\in G_{1}$ for each $i\in\{1,2,\cdots,l\}$, then there exists a random subsequence $\{(x_{1}^{M_{n}},x_{2}^{M_{n}},\cdots,x_{l}^{M_{n}}),n\in\mathbb{N}\}$ of $\{(x_{1}^{M_{n}^{(0)}},x_{2}^{M_{n}^{(0)}},\cdots,x_{l}^{M_{n}^{(0)}}),n\in\mathbb{N}\}$ such that $\{x_{i}^{M_{n}},n\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y_{i}$ and $\{f(x_{i}^{M_{n}}),n\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(y_{i})$ for each $i\in\{1,2,\cdots,l\}$.
Lemma 4.10 below is the $L^{0}$-simplex version of the random Brouwer fixed point theorem in $RN$ modules.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $RN$ module $E$ and $f:S\rightarrow S$ be a $\sigma$-stable and random sequentially continuous mapping. Then $f$ has a fixed point.
@@
Let $\mathscr{S}$ be an $L^{0}$-barycentric subdivision of $S$. For $\phi$ and $\mathscr{S}$, by Theorem 2.9 there exists an (n1)(n-1)-dimensional $L^{0}$-simplex $S^{1}:=Conv_{L^{0}}(\{y_{1}^{1},y_{2}^{1},\cdots,y_{n}^{1}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}$, namely, there exist a finite family $\{S_{k}^{1},k=1\sim M\}$ of (n1)(n-1)-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{k}^{1},k=1\sim M\}\in p(1)$ such that
@@
Let $\mathscr{S}_{1}$ be an $L^{0}$-barycentric subdivision of $S^{1}$. Now, similarly, for $\phi$ and $\mathscr{S}_{1}$, by Theorem 2.9 there exists an (n1)(n-1)-dimensional $L^{0}$-simplex $S^{2}:=Conv_{L^{0}}(\{y_{1}^{2},y_{2}^{2},\cdots,y_{n}^{2}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{1}$, such that $S^{2}$ is completely labeled by $\phi$ and satisfies that
@@
By induction, for each $m\in\mathbb{N}$, there exists an (n1)(n-1)-dimensional $L^{0}$-simplex $S^{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\})$, which is a finite concatenation of some finite family of $L^{0}$-simplexes in $\mathscr{S}_{m-1}$, such that $S^{m}$ is completely labeled by $\phi$ and satisfies that
@@
Then $\{(y_{1}^{m},y_{2}^{m}\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ is a sequence in $S^{(n)}$, where $S^{(n)}$ is the $n$-th Cartesian power set of $S$.
For $\{y_{1}^{m},m\in\mathbb{N}\}$, since $S$ is random sequentially compact by Remark 4.4, there exists a random subsequence $\{y_{1}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converging in $\mathcal{T}_{\varepsilon,\lambda}$ to some $y$ in $S$. Since $\{diam(S^{m}),m\in\mathbb{N}\}$ converges $a.s.$ to 0, by [53, Lemma 2.11] $\{d^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges $a.s.$ (of course, also in $\mathcal{T}_{\varepsilon,\lambda}$) to 0, where
@@
implying that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$. To sum up, $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}})$, $k\in\mathbb{N}\}$ is a random subsequence of $\{(y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}),m\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}^{\prime}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ for any $i\in\{1,2,\cdots,n\}$. Since $f$ is $\sigma$-stable and random sequentially continuous, by Proposition 4.9 there exists a random subsequence $\{(y_{1}^{m_{k}},y_{2}^{m_{k}},\cdots,y_{n}^{m_{k}}),k\in\mathbb{N}\}$ of $\{(y_{1}^{m_{k}^{\prime}},y_{2}^{m_{k}^{\prime}},\cdots,y_{n}^{m_{k}^{\prime}}),k\in\mathbb{N}\}$ such that $\{y_{i}^{m_{k}},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $y$ and $\{f(y_{i}^{m_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(y)$ for each $i\in\{1,2,\cdots,n\}$.
Let $y=\sum_{i=1}^{n}\lambda_{i}x_{i}$ and $f(y)=\sum_{i=1}^{n}\mu_{i}x_{i}$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ are such that $\sum_{i=1}^{n}\lambda_{i}=\sum_{i=1}^{n}\mu_{i}=1$. Next, we will show that $f(y)=y$, namely, $y$ is a fixed point of $f$.
@@
$=j.$
So $\lambda_{j}^{m_{k},j}\geq\mu_{j}^{m_{k},j}$ by Lemma 4.8.
Furthermore, by [27, Lemma 5.5], $\{\lambda_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\lambda_{j}$ and $\{\mu_{j}^{m_{k},j},k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $\mu_{j}$ for each $j\in\{1,2,\cdots,n\}$, then
@@
Lemma 4.10 is a generalization of [12, Theorem 2.3], since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is an $RN$ module, and a $\sigma$-stable a.s. sequentially continuous mapping is, obviously, random sequentially continuous, as noted in Remark 4.6. Since random sequential continuity is much more complicated than a.s. sequential continuity, Proposition 4.9 has played a key role in the treatment of random sequential continuity in the proof of Lemma 4.10. Besides, we would also like to emphasize another difference between the proofs of Lemma 4.10 and [12, Theorem 2.3]: the sequence $\{S^{m},m\in\mathbb{N}\}$ of (n1)(n-1)-dimensional $L^{0}$-simplexes, which is constructed in Lemma 4.10, has the nice property that $S^{m+1}\subseteq S^{m}$ for each $m\in\mathbb{N}$. Contrarily, the sequence $\{S^{m},m\in\mathbb{N}\}$ constructed in [12, Theorem 2.3] does not have this property. The nice property of our sequence will be used to study other problems in the future. Besides, the proof of Lemma 4.10 is simpler and clearer than that of [12, Theorem 2.3] since we only employ a single proper $L^{0}$-labeling function throughout the proof of Lemma 4.10.
Let $H$ be a Hilbert space and $G$ be a nonempty closed convex subset of $H$. Then it is well known that the classical projection $P_{G}$ of $H$ onto $G$ is nonexpensive. By a completely similar method, one can prove the following lemma.
@@
Let $(H,\langle\cdot,\cdot\rangle)$ be a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ and $G$ be a nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $H$. Then for any $x\in H$, there exists an unique element $P_{G}(x)$ in $G$ such that
@@
where $\|\cdot\|$ is the $L^{0}$-norm induced by the $L^{0}$-inner product $\langle\cdot,\cdot\rangle$. Furthermore, this defines a mapping $P_{G}:H\rightarrow G$, called the random projection onto $G$ that is nonexpensive, namely,
@@
In Lemma 4.12, since $G$ is $\sigma$-stable and $P_{G}$ is $L^{0}$-Lipschitzian, it follows that $P_{G}$ is $\sigma$-stable by [28, Lemma 2.11].
@@
Since $G$ is $a.s.$ bounded, there exists an $L^{0}$-simplex $S$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ such that $G\subseteq S$. Since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module, let $h:S\rightarrow G$ be the restriction of $P_{G}$ to $S$, where $P_{G}$ is the random projection of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ onto $G$. It is clear that $h$ is $\mathcal{T}_{c}$-continuous.
@@
Proposition 4.13 below was first established as [27, Lemma 4.7], which is merely a reformulation of Theorem 2.1 in the case of a finitely generated $RN$ module. For the proof of Proposition 4.13, we refer the reader to [27].
@@
5 The equivalence between Proposition 2.3 and Proposition 2.4
The main purpose of this section is to establish the equivalence between Proposition 2.3 and Proposition 2.4. In fact, that Proposition 2.3 implies Proposition 2.4 is obvious, but the reverse implication is another matter. Owing to the profound result of Ponosov [44], which is the solution to a Nemytskij conjecture (see also Lemma 5.2 below), we can prove the reverse implication. It is in proving the reverse implication that we also obtain the following two somewhat surprising by-products — Theorem 5.6 and Proposition 5.8. Theorem 5.6 shows that both Proposition 2.3 and Proposition 2.4 are equivalent to the random version of the original Brouwer fixed point theorem (namely, part (2) of Theorem 5.6, whose proof only needs the use of the measurable selection theorem — [55, Theorem 4.2]). Proposition 5.8 shows that the $\mathcal{T}_{\varepsilon,\lambda}$-continuity and a.s. sequential continuity coincide for a $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$.
@@
Sufficiency. For any $a\in B_{\mathcal{F}}$ and any $x,y$ in $G$, from the $\sigma$-stability of $h$ one has that $h(I_{a}x+I_{a^{c}}y)=I_{a}h(x)+I_{a^{c}}h(y)$. Then, when $I_{a}x=I_{a}y$, $h(y)=I_{a}h(x)+I_{a^{c}}h(y)$, so $I_{a}h(x)=I_{a}h(y)$ holds, which amounts to saying that $h$ is local. ∎
Lemma 5.2 below is a special case of the main result of [44], which is sufficient for our purpose.
@@
Let $X$ and $Y$ be two complete separable metric spaces, $(\Omega,\mathcal{F},P)$ a complete probability space and $h:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ a local and continuous-in-probability operator. Then there exists a Carathéodory function $f:\Omega\times X\rightarrow Y$ (namely, $f(\cdot,x)$ is measurable for any $x\in X$ and $f(\omega,\cdot)$ is continuous for any $\omega\in\Omega$) such that $h=h_{f}$. Here, $h_{f}:L^{0}(\mathcal{F},X)\rightarrow L^{0}(\mathcal{F},Y)$ is defined by $h_{f}(x)=$ the equivalence class of $f(\cdot,x^{0}(\cdot))$, where $x^{0}$ is an arbitrarily chosen representative of $x$.
In fact, Lemma 5.2 has implied that $h$ is a.s. sequentially continuous, namely, $\{h(x_{n}),n\in\mathbb{N}\}$ converges a.s. to $h(x)$ whenever $\{x_{n},n\in\mathbb{N}\}$ converges a.s. to $x$.